Bob earned 420 on his investment in bonds and stocks. If his bonds return 2% and his stocks returned 6% and his total investment was 10,000 how much did he invest in
Answer by mananth(16946)

Bonds 2.00% per annum ------------- Amount invested =x
Stocks 6.00% per annum ------------ Amount invested = y
10000
Interest----- $420

Bonds 2.00% per annum ---x
Stocks 6.00% per annum ---y
Total investment
x + 1 y= 10000 -------------1
Interest on both investments
2.00% x + 6.00% y= 420
Multiply by 2
2 x + 6 y= 42000.00 --------2
Multiply (1) by -2
we get
-2 x -2 y= -20000.00
Add this to (2)
0 x 4 y= 22000
divide by 4
y = 5500
Bonds 2.00% $ 4500
Stocks 6.00% $ 5500

CHECK
4500 --------- 2.00% ------- 90.00
5500 ------------- 6.00% ------- 330.00
Total -------------------- 420.00

m.ananth@hotmail.ca